# Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Transfer in Sense-Aware Tasks

### v1

@@ 4.3 Lexical Semantic Change Detection
We evaluate XL-LEXEME, trained on multilingual and interlingual data (all WiC datasets plus AM2iCO), against the same mono models from earlier analyses. As in WiC, multi is matched or outperformed in all but two languages by mono English or German. XL-LEXEME outperforms only on German, but all well-trained mono models perform well on this task. Results correlate strongly with pretraining sizes in all three models (Figure 2), echoing WiC patterns.
Zero-shot transfer performs on par with full-shot. On German, mono French and English outperform the German model, and English is most effective on Spanish, Chinese, and Norwegian. PRT confirms these trends, with English showing the highest correlation. Across mBERT and BLOOM, mono models often match or outperform multi models (Tables 24-26).666LLaMA models were not evaluated on LSCD as prompting generative models with hundreds of usages to produce a scalar ‘change score’ is not practically feasible.
@@ 5 Discussion
Quantity, not Multilinguality
The multi model only partly outperformed monolingual models, with the english model often matching or surpassing it, even under the full f.t. condition. Strict control conditions across tests show that multilinguality does not inherently improve zero-shot transfer. Instead, its performance boost is due to factors like fine-tuning size and full-shot conditions. These results challenge the assumption that training on multilingual data inherently aids transfer; it is the quantity of training examples, not their multilinguality, that improves performance.
Quality, not only Quantity
mono English frequently outperformed multi model on most WiC and LSCD tasks, suggesting a special advantage. These results hold not only for fixed f.t, where quantity is equal, but even for full f.t, where english has far less training data compared to other languages. We suggest this advantage reflects the higher quality of English WiC, likely due to differences in dataset curation, annotation protocols, sense granularity, and text quality, echoing the conclusions from Philippy et al. (2023)’s survey.
Underlying Factors of Successful Transfer
@@
The strong performance on English tasks observed in our results is likely due to its prevalence during pretraining, and supported by the strong correlations with PT-size reported in §4.2. The contrasting results from MuRIL, where transfer was strongest to Hindi – when it had the PT-size advantage over English – further underscore this point, providing an informative exception.
@@
The influence of pretraining corpus size on cross-lingual transfer has been previously established Lauscher et al. (2020); Srinivasan et al. (2021); Ahuja et al. (2023), as well as the correlation between language similarity and transfer performance Wu and Dredze (2019); Pires et al. (2019); K et al. (2020). However, these factors affect tasks differently: pretraining size correlates more with semantic tasks (e.g., NLI, QA), while linguistic similarity benefits syntactic tasks (e.g., POS tagging, dependency parsing). This contrast is supported by results from Lauscher et al. (2020) and echoes Philippy et al. (2023) who argued that transfer is shaped by a combination of interacting factors.
@@
This distinction sheds light on the types of linguistic knowledge that facilitate transfer. We propose that syntactic tasks rely more heavily on structural similarities across languages, and thus benefit from linguistic proximity. As syntactic patterns diverge across typologically distant languages, their utility in transfer diminishes. In contrast, semantic tasks - particularly those involving complex reasoning - are more dependent on the diversity and scale of pretraining data, which exposes models to a wider range of meaning representations. We suggest that sense-aware tasks, such as polysemy disambiguation and semantic change, follow this semantic pattern. Given the language-specific nature of word senses, syntactic similarity offers limited benefit of knowledge to transfer. Our findings confirm this: successful transfer in these tasks depends more on the size of multilingual pretraining than on linguistic similarity.
When More is Less
We include LLaMA in our analysis as a representative of modern generative large language models. It has more than 14 times more parameters than XLM-R, and was pretrained on orders of magnitude more data, likely including data very similar to our own Sainz et al. (2023). Nevertheless, LLaMA consistently underperforms relative to XLM-R across the majority of languages. This further supports our conclusion that training on more data and languages, even with the most modern models, does not guarantee improved cross-lingual transfer.
@@ 6 Conclusion
While further research is needed to determine whether these patterns generalize across other tasks, they may underline broader implications for multilingual NLP research that traditionally favored multilingual quantity over data quality. As expanding datasets across multiple languages is costly, and sometimes not practical for some tasks and languages, future research should prioritize understanding the characteristics of high-quality datasets, exposing potential confounds, and optimizing training resources for effective cross-lingual transfer.
@@ 7 Limitations
Zero-shot transfer enables models to perform tasks in languages without task-specific training data. However, its effectiveness depends on the pretrained model’s ability to capture the linguistic properties of both the source and target languages, as well as the availability of sufficient, high-quality data in the fine-tuning language. Our study is constrained by this limitation, as we were only able to train models on a limited set of languages with highly imbalanced data distributions. As in broader low-resource language research, high-resource languages dominate, while lower-resource languages often lack sufficient data for effective fine-tuning.
@@
Our evaluation of LLaMA is limited to the WiC task, and several factors constrain its interpretability. First, its pretraining data is unknown so we could not assess the correlation of performance with pretraining size. Second, LLaMA’s pretraining could include the datasets used in our training or evaluation. And lastly, adapting WiC to LLaMA’s generative format (see Appendix I.2) introduced further limitations, such as the lack of a threshold to control sense discrimination, and a cue about the task for LLaMA, which encoder-based models lacked. Considering these, LLaMA is not directly comparable to the other models. Future work is needed to better understand these architectural differences.
Despite the exclusion of most LLMs, our study remains broadly relevant. Simpler encoder-based models continue to play a key role, especially in scenarios requiring efficient, resource-conscious deployment. Unlike LLMs, they can run on edge devices, perform better in low-resource settings and specialized domains, and are far more cost-effective. As LLMs remain computationally expensive and inaccessible in many real-world applications, encoder models will continue to be important for both research and industry.
@@ 9 Acknowledgments
We are grateful to Francesco Periti and Dominik Schlechtweg for their invaluable feedback and insightful advice on this paper.
@@ Appendix D Zero-Shot Comparison of mBERT and BLOOM Models
Appendix E Error Analysis
We perform error analysis to further analyse how, and if multilingual models differ from their monolingual counterparts.
@@ Appendix E Error Analysis
Appendix F Pretraining Sizes of Languages Used in Our Analysis
@@
Language proportions for LLaMA-3–8B-Instruct’s pretraining corpus are not publicly available.
Appendix G All Results on Monolingual and Cross-Lingual WiC tasks
@@
Appendix H Performance of XLM-R, mBERT and BLOOM-Based Models on LSCD Tasks
@@
Appendix I LLaMA
I.1 LLaMA outperformed by XLM-R
@@
While we include LLaMA in our evaluation to provide a comparison with recent large language models and SOTA instruction-tuned architectures, our results suggest that encoder-based models like XLM-R remain more effective for in-context polysemy disambiguation. Despite the generative capabilities of LLaMA, it underperforms in most cases compared to XLM-R, particularly when ample training data is available. This highlights the continued relevance of encoder-based models, which demonstrate stronger task-specific performance and greater cross-lingual transfer.
I.2 Prompts and Response Parsing
System message:
User message:
if "1" in raw and "0" not in raw:
return "1"
elif "0" in raw and "1" not in raw:
return "0"
else:
return None # counts as incorrect
Appendix J LLaMA Fine-Tuning Parameters

### v2

@@ 4.3 Lexical Semantic Change Detection
We evaluate XL-LEXEME (Cassotti et al., 2023), trained on multilingual and inter-lingual data (all WiC datasets plus AM2iCO), against the same mono models from earlier analyses. As in WiC, multi is matched or outperformed in all but two languages by mono English or German. XL-LEXEME outperforms only on German, but all well-trained mono models perform well on this task. Notably, mono English outperforms all other models on average and in 4 out of 8 test languages. Results correlate strongly with pretraining sizes in all three models (Figure 2), echoing WiC patterns.
Zero-shot transfer performs on par with full-shot. On German, mono French and English outperform the German model, and English is most effective on Spanish, Chinese, and Norwegian. PRT confirms these trends, with English showing the highest correlation. Across mBERT and BLOOM, mono models often match or outperform multi models (Tables 26-28).777LLaMA models were not evaluated on LSCD as prompting generative models with hundreds of usages to produce a scalar ‘change score’ is not practically feasible.
@@ 5 Discussion
Disentangling the effects of multilinguality from confounding factors such as data quality and pretraining exposure is inherently challenging. Rather than providing a single definitive explanation, we identify consistent trends and highlight effects that reflect the nuanced nature of cross-lingual transfer.
Quantity over Multilinguality
Our results show that the amount of fine-tuning data matters more than the number of languages included during training, as full f.t. mono models trained on more data typically outperform the fixed f.t. multi model. In both fine-tuning conditions, the multi model only partially outperforms monolingual models, with the english model often matching or exceeding it, even in the full f.t. setting. Taken together, this pattern of results challenges the assumption that multilingual training inherently improves zero-shot transfer; rather, data quantity drives most gains. Notably, Berend (2022), who also tested the role of multilinguality in word sense disambiguation but in the pretraining stage, reached the same conclusions. Our results not only complement those of Berend but are also more practical as multilingual models are in much wider use than monolingual ones in transfer scenarios, where the latter are actually lacking for many languages.
Quality Beyond Quantity
The frequent outperformance of the mono English model on most tasks could point to data quality as an important factor. Certainly, the quality of training data can impact performance, and consequently, transfer to other languages. English may enjoy datasets of higher quality, often attributed to their curation, annotation protocols, sense granularity and text quality Philippy et al. (2023). However, assessing or normalizing data quality across languages for polysemy disambiguation is nearly impossible without costly native-speaker annotations. Thus, this factor remains difficult to study directly.
Other Underlying Factors
@@
english’s strong performance may be tied to its dominance in pretraining data which is supported by its strong overall correlations reported in §4.2. The contrasting results from MuRIL, where transfer was strongest to Hindi when Hindi had a pretraining size advantage over English, further underscore this point and provide an informative exception.
@@
Our results show that both pretraining size and language similarity are tied to transfer, but the former has a stronger correlation. The influence of pretraining size on cross-lingual transfer is well established Lauscher et al. (2020); Srinivasan et al. (2021); Ahuja et al. (2023), as the correlation between language similarity and transfer performance Wu and Dredze (2019); Pires et al. (2019); K et al. (2020). However, while both variables clearly influence transfer, the effect seems to depend on the specific task on which transfer is evaluated: pretraining size correlates more with semantic tasks (e.g., NLI, QA), while linguistic similarity benefits syntactic tasks (e.g., POS tagging, dependency parsing) Lauscher et al. (2020); Philippy et al. (2023).
@@
We suggest that sense-aware tasks, such as polysemy disambiguation and semantic change, should follow the semantic route. Given the language-specific nature of word senses, syntactic similarity offers limited benefit in the transfer of knowledge. Our findings confirm this: transfer performance in these tasks depends more on pretraining size than on linguistic similarity. The rationale of this interpretation is rooted in the type and relevance of linguistic knowledge that is being transferred Rajaee and Monz (2024); Goldman et al. (2025). We propose that syntactic tasks rely more heavily on structural similarities across languages, and thus benefit from linguistic similarity. As syntactic patterns diverge across typologically distant languages, their utility in transfer diminishes. In contrast, semantic tasks are more dependent on the diversity and scale of pretraining data, which exposes models to a wider range of meaning representations.
Lexical Semantic Change
The results on LSCD align with the conclusions observed in the WiC tasks: multilinguality does not necessarily lead to improved performance, and monolingual models, particularly english, often outperform multilingual ones. This holds even when the test language is included in the multilingual setup (and, of course, absent from the monolingual model), as seen in the cases of Spanish, Chinese, and Norwegian, where the English model outperformed both multilingual models. These findings reinforce the paper’s central claim: multilinguality is neither necessary nor inherently beneficial for effective transfer, and monolingual models can surpass multilingual ones even under comparison conditions that favor the latter. In such cases, the multilingual model benefits from full-shot exposure to the test language, while the monolingual model is evaluated in a zero-shot setting—yet still performs better. These results further emphasize the role of dataset quality in transfer, suggesting that the perceived advantages of multilinguality may be confounded by differences in training data quality.
As lexical semantic change gains popularity and more languages are being studied using computational methods, it is important not to perpetuate the misconception that multilingual models are always the best solution—not even when the target language is included in the multilingual training setup. This conclusion is further supported by Baes et al. (2025), who recently showed that the state-of-the-art model for LSCD—the multilingual XL-LEXEME by Cassotti et al. (2023) also included in our comparisons—is not the best choice across different semantic change scenarios. Their findings call for a more nuanced approach to model selection—one that accounts for the specific conditions of the linguistic inquiry being studied, whether it is a different language or type of change.
Large Models and Architectures
Notably, larger models like LLaMA, pretrained on much more data, underperform compared to smaller models like XLM-R. This corroborates our main finding that increasing the amount of training data alone – whether in terms of language coverage or volume – does not guarantee better transfer. However, architectural differences (encoder vs. decoder), hyperparameter tuning, and adaptation methods likely also influence performance and require further investigation. Additionally, Berend (2022) suggests that multilingual models may not even be necessary for effective transfer, highlighting that monolingual pretrained models can achieve strong cross-lingual generalization on sense-aware tasks when combined with appropriate adaptation techniques.
@@ 6 Conclusion
While further research is needed to determine whether these patterns generalize across other tasks, they may underline broader implications for multilingual NLP research that traditionally favored multilingual quantity over data quality. As expanding datasets across multiple languages is costly, and sometimes not practical for some tasks and languages, future research should prioritize understanding the characteristics of training regimes that facilitate transfer in order to optimize training resources for effective cross-lingual transfer.
@@ 7 Limitations
Zero-shot transfer enables models to perform tasks in languages without task-specific training data. However, its effectiveness depends on the pretrained model’s ability to capture the linguistic properties of both the source and target languages, as well as the availability of sufficient, high-quality data in the fine-tuning language. Our study is constrained by this limitation, as we were only able to train models on a limited set of languages with highly imbalanced data distributions. As in broader cross-lingual transfer research, high-resource languages dominate, while lower-resource languages often lack sufficient data for effective fine-tuning.
@@
Our evaluation of LLaMA is limited to the WiC task, and several factors constrain its interpretability. First, its pretraining data is unknown so we could not assess the correlation of performance with pretraining size. Second, LLaMA’s pretraining could include the datasets used in our training or evaluation. And lastly, adapting WiC to LLaMA’s generative format (see Appendix L.1) introduced further limitations, such as the lack of a threshold to control sense discrimination, and a cue about the task for LLaMA, which encoder-based models lacked. Considering these, LLaMA is not directly comparable to the other models. Future work is needed to better understand the impact of these architectural differences. Due to computational resource constraints, we were unable to include other larger-scale models (8B+ parameters) or mixture-of-experts (MoE) variants in our experiments. Despite this, our study remains broadly relevant as encoder-based models remain crucial for tasks like classification, regression, and ranking in retrieval systems, where their efficiency and lower resource demands make them well-suited. Additionally, on-device performance requires fast inference and low memory usage. While recent hardware and quantization techniques enable some larger models to run locally on powerful devices, encoder models remain preferred for many real-world applications that require fast, cost-effective, and scalable solutions, especially in specialized or low-resource settings.
@@ 9 Acknowledgments
We are grateful to Francesco Periti and Dominik Schlechtweg for their invaluable feedback and insightful advice on this manuscript, and to Mahmud Akhter and Pierluigi Cassotti for their support.
@@ Appendix D Zero-Shot Comparison of mBERT and BLOOM Models
Appendix E On the Lack of Extensive Hyperparameter Tuning
To keep cross-model and cross-language comparisons fair and interpretable, we deliberately avoided extensive per-model or per-language hyperparameter tuning. Instead, we adopted a single configuration taken from independent prior work and applied it uniformly across all models and settings, with an identical early-stopping rule. This choice limits researcher degrees of freedom and reduces the risk of inadvertently “tuning into” our research hypothesis, which can happen when many parameters are adjusted differently across conditions. While more aggressive tuning could raise the peak performance of individual systems, it would blur causal attribution and undermine the comparability that our study seeks to emphasize.
Appendix F Error Analysis
We perform error analysis to further analyze how, and if multilingual models differ from their monolingual counterparts.
@@ Appendix F Error Analysis
Appendix G Pretraining Sizes of Languages Used in Our Analysis
@@
Language proportions for LLaMA-3–8B-Instruct’s pretraining corpus are not publicly available.77footnotetext: The pretraining corpus sizes for languages used to train mBERT are also not publicly reported. For these, we rely on averaged estimates from Wu and Dredze (2020), who measured the Wikipedia size for each language included in mBERT’s training data.
Appendix H All Results on Monolingual and Cross-Lingual WiC tasks
@@
Appendix I Performance of Pretrained Models without Fine-Tuning
It is logical to ask whether pretrained multilingual language models already separate word senses in a no-fine-tuning setting. Because contextual embeddings integrate sentence context via attention, one might expect embeddings of mole in “burrowing mammal” vs. “skin blemish” to diverge more than two occurrences of the same sense, even without fine-tuning. To probe this, we evaluate pretrained models without task-specific training by classifying sentence pairs using cosine distance between target-word embeddings, consistent with our main protocol.
Threshold selection and considerations.
In our main experiment setup, each fine-tuned encoder uses a cosine-distance decision threshold set on dev data from its fine-tuning language (or multilingual dev for MULTI models). For encoders without fine-tuning, no model-specific dev data exists, leaving only flawed alternatives: (A) calibrate once on pooled multilingual dev data (introducing mixture bias), (B) tune separately on each test set (test leakage, inflated scores), or (C) calibrate per language (uses target-language supervision unavailable to the pretrained encoders). We adopt Option A as the least biased feasible choice. We sweep candidate thresholds on the pooled dev data to maximize accuracy and then fix that single global threshold for each encoder across all datasets. This avoids test leakage and keeps calibration constant, but the global threshold inevitably reflects mixture statistics and tailors to the average of all the data, and can favor high-resource languages. Accordingly, these numbers are a sanity check for learning rather than directly comparable performance estimates.
Results.
Figure 7 shows that XLM-R, mBERT, and BLOOM consistently benefit from fine-tuning across nearly all conditions, with the few exceptions concentrated in low-data fine-tuning settings (e.g., Chinese or Fixed F.T.) when evaluated on Hindi. These gains indicate that fine-tuning substantially improves sense discrimination beyond what is present in the pretrained geometry.
For LLaMA, the dynamics differ slightly. The generative model is prompted (see Figure 9 for the prompt) rather than based on a cosine-distance threshold. Its zero-shot instruction-following yields a stronger baseline than the pretrained encoders on most tasks; accordingly, despite using the same WiC training data, its absolute gains from fine-tuning are smaller. This reflects an architectural/interface contrast—encoders benefit from shaping the embedding space and a calibrated decision threshold, whereas LLaMA already executes the task from the prompt and changes less with additional supervision.
We include full results for the no-fine-tuning baselines in Table 24. These baselines show that sense discrimination improves with fine-tuning; however, they are not intended for head-to-head comparison with fine-tuned models due to the necessarily different threshold calibration.
Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI XLM-R 62.2 62.2 61.6 58.7 58.0 58.5 60.6 56.1 56.1 53.1 53.9 53.1 57.0 54.8 56.9 55.7 58.7 60.0 66.1 mBERT 56.2 58.5 51.2 52.1 53.7 58.1 54.2 54.4 52.8 54.7 55.2 52.6 54.7 50.3 56.4 56.8 62.6 53.1 63.8 BLOOM 61.1 50.0 50.0 50.0 50.0 49.9 50.0 49.0 50.0 50.1 49.9 50.0 50.1 49.7 50.3 50.0 50.0 53.9 63.6 LLaMA 67.2 64.2 65.4 58.0 62.8 55.4 63.2 56.9 55.7 58.4 57.9 53.8 53.4 52.8 46.5 54.1 53.7 70.6 62.5 MuRIL 50.0 49.9 57.4 51.8 50.4 49.8 49.9 48.5 50.8 50.6 56.1 58.2 49.7 50.0 56.2 50.2 53.2 50.4 59.8
Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH XLM-R 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 49.9 50.1 50.0 50.0 mBERT 50.0 50.0 50.1 50.1 50.0 50.0 50.0 50.0 50.0 50.0 49.9 49.9 50.0 50.0 BLOOM 50.0 50.0 50.0 50.8 49.9 50.1 50.2 50.1 49.8 49.9 49.7 49.9 50.0 49.8 LLaMA 58.3 55.7 61.6 55.3 57.3 59.3 58.7 54.7 53.2 57.9 57.7 56.8 59.0 60.3 MuRIL 50.5 49.7 50.3 50.0 49.8 50.6 49.7 50.6 50.0 49.6 49.8 50.1 52.2 50.0
Appendix J Performance of XLM-R, mBERT and BLOOM-Based Models on LSCD Tasks
@@ Results.
Appendix K LLaMA
K.1 LLaMA outperformed by XLM-R
@@
While we include LLaMA in our evaluation to provide a comparison with recent large language models (LLMs) and instruction-tuned architectures, our results suggest that encoder-based models like XLM-R remain more effective for in-context polysemy disambiguation. Despite the generative capabilities of LLaMA, it underperforms on these tasks in most languages compared to XLM-R, particularly when ample training data is available. This highlights the continued relevance of encoder-based models, which demonstrate stronger task-specific performance and greater cross-lingual transfer.
Appendix L LLaMA Fine-Tuning Parameters
@@
L.1 Prompts and Response Parsing
System message:
User message:
if "1" in output and "0" not in output:
return "1"
elif "0" in output and "1" not in output:
return "0"
else:
return None # counts as incorrect